If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?: If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by ptaylor(1184)

Let x=amount of time it will it take all 3 of them working together to mix the 20 drinks. {Our final equation will be of the form: Rate at which they mix drinks when working together (drinks/min) times x minutes equals 20 drinks}
Steven mixes at the rate of 4 drinks per minute
Sue mixes at the rate of 2 drinks per minute
Jack mixes at the rate of 20/15 or 4/3=1 1/3 drinks per minute
Together, they mix at the rate of 4 + 2 + (1 1/3)=7 1/3 drinks per minute
So, our equation to solve is:
(7 1/3)*x=20 divide each side by 7 1/3
x=2.73 minutes---------------------amount of time it will it take all 3 of them working together to mix the 20 drinks
CK
2.73*7.33=20
20.019~~~~20


Hope this helps---ptaylor


Question 149044: If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?: If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by Electrified Levi(31)

Hi, Hope I can help,
.
If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
.
First, you have to make sure that all the measurements are the same(if it says Bob can do it in 5 hours, and John can do it in 45 minutes, you have to convert minutes to hours, or hours to minutes) They are all the same in our problem
.
This is the way I usually solve these types of problems
.
You are trying to find out how long it will take them, doing the job together.
.
Here is the formula I use
.
 (x/(how long it takes for person(1)))+(x/(how long it takes for person(2)))+(x/(how long it takes for person(3))) = 1 ( If there was another you would add
.
(x/(how long it takes for person(4))) , and so on)
.
The "x" is how long it will take for all of them together to get the job done
.
If you add the fractions together it will equal 1 job
.
We can now replace the bottom numbers(denominators) with "5","10","15"( that's how long it takes each person)
.
 (x/5)+(x/10)+(x/15) = 1
.
We can now solve for "x", we will multiply everything by "30" to get rid of the denominators, and fractions
.
 (30)(x/5)+(30)(x/10)+(30)(x/15) = (30)1
.
It will become
.
 6x + 3x + 2x = 30
.
We will add the left side
.
 11x = 30
.
We can divide each side by "11"
.
 11x/11 = 30/11
.
 x = 30/11
.
If they all work together they can get the job done in 2  8/11 minutes, we can check by replacing "x" with "2  8/11 " or " 30/11 " in our equation
.
 ((30/11)/5)+((30/11)/10)+((30/11)/15) = 1
.
 (6/11)+(3/11)+(2/11) = 1
.
 11/11 = 1
.
 1 = 1
.
They can get the job done in  30/11 minutes, or 2  8/11 minutes
.
Hope I helped, Levi